### v1

# Learning to See and Act: Task-Aware View Planning for Robotic Manipulation

Abstract
Recent vision-language-action (VLA) models for multi-task robotic manipulation commonly rely on static viewpoints and shared visual encoders, which limit 3D perception and cause task interference, hindering robustness and generalization. In this work, we propose Task-Aware View Planning (TAVP), a framework designed to overcome these challenges by integrating active view planning with task-specific representation learning. TAVP employs an efficient exploration policy, accelerated by a novel pseudo-environment, to actively acquire informative views. Furthermore, we introduce a Mixture-of-Experts (MoE) visual encoder to disentangle features across different tasks, boosting both representation fidelity and task generalization. By learning to see the world in a task-aware way, TAVP generates more complete and discriminative visual representations, demonstrating significantly enhanced action prediction across a wide array of manipulation challenges. Extensive experiments on RLBench tasks show that our proposed TAVP model achieves superior performance over state-of-the-art fixed-view approaches. Visual results and code are provided at: TAVP.
Introduction
A general-purpose robotic system is expected to accurately perform a wide range of manipulation tasks [7]. With the rapid development of vision-language-action (VLA) models [39, 119, 73, 94, 74, 111, 84, 115, 116, 117, 38, 118] and transformer-based architectures [122, 119, 73, 31, 30], many end-to-end robotic manipulation frameworks (e.g., OpenVLA[39], $\pi_{0}$[120]) have achieved promising results on complex and fine-grained tasks. Typically, these models rely on observations captured from a single view [39, 30, 42] or a few fixed viewpoints [158, 157] to guide action prediction. While this may suffice in simple environments with only a few objects, it becomes problematic in cluttered or dynamic scenes—particularly when objects or the end-effector are in motion during task execution.
In such scenarios, fixed viewpoints often lead to occlusions of the target object or the end-effector, resulting in incomplete scene understanding and inaccurate action predictions. As shown in Fig. 1, given the instruction “Put the sugar in the cupboard”, images captured from three fixed cameras either miss the cupboard or fail to include the end-effector grasped with sugar. Such incomplete visual observations prevent the robot from correctly interpreting the scene, ultimately leading to suboptimal or failed actions.
Figure 1: Motivation Illustration. Observations captured from fixed cameras often miss parts of the target objects. For example, the front view only captures the cupboard (highlighted with a red circle), while the left and right shoulder views only show the sugar (already grasped by the end-effector and highlighted with green circles). These incomplete observations may lead to failed operations. In contrast, our proposed TAVP is designed to dynamically explore and re-render informative viewpoints that maximize coverage of target-relevant information, thereby improving the reliability of manipulation outcomes.
Beyond acquiring complete and informative visual features, a scalable and task-sensitive visual encoder is also critical for accurate perception in multi-task settings. Although prior language-conditioned transformer-based methods, such as RVT [155] and RVT-2 [143], enable multi-task capabilities by attending to instruction tokens, their shared encoders still suffer from task interference, especially when handling visually and semantically diverse tasks (e.g., pick an apple vs. open a drawer). This entanglement limits the generalization and scalability of such models across a broad range of manipulation behaviors.
To address these challenges, we propose the Task-Aware View Planning (TAVP) model, which integrates viewpoint exploration with task-aware visual feature extraction for robust and generalizable multi-task robotic manipulation. TAVP actively explores informative viewpoints using a delicately designed efficient reinforcement learning–based exploration policy, while employing a scalable task-aware Mixture-of-Experts (MoE) visual encoder to extract task-specific features. This design enables the model to obtain more complete and discriminative visual information, thereby improving robustness and generalization in multi-task action prediction.
Specifically, our task-aware Mixture-of-Experts module, termed TaskMoE, improves multi-task accuracy and generalization through two key designs: (1) a dynamic expert routing mechanism guided by fused instruction and scene cues, rather than relying solely on task identifiers; and (2) a decoupled gating strategy that enables parameter sharing among semantically similar tasks, while isolating routing for semantically diverse ones. Based on the extracted task-specific features, we introduce the Multi-Viewpoint Exploration Policy (MVEP) to train a viewpoint exploration policy that selects optimal camera poses for each task. This Policy predicts the optimal camera re-rendering viewpoints based on the global point cloud. It then renders corresponding 2D images from these viewpoints for subsequent feature extraction and action prediction. These features possess enhanced global awareness with richer information, capable of improving robotic performance.
To comprehensively evaluate the effectiveness of our TAVP in enhancing robotic spatial understanding and action generation performance, as well as its multi-task perception and decision-making capabilities, we conducted rigorous benchmark testing on 18 diverse RLBench [19] simulation tasks covering multiple tasks and scenarios. Our contributions are summarized as follows:
  • Multi-Viewpoint Exploration Policy (MVEP) : The MVEP enables robots to actively explore optimal observation viewpoints (dynamic multi-view re-rendering), effectively addressing occlusion and insufficient viewpoint problems, and enhancing 3D perception capabilities.
  • Task-aware Mixture-of-Experts (TaskMoE): We introduce TaskMoE, which dynamically selects perception and action generation experts according to the task and its related instruction and scene visual information, improving multi-task processing capability and robustness.
  • Superior Robotic Manipulation Performance: Extensive experiments on 18 RLBench simulation tasks demonstrate that our TAVP significantly outperforms existing baselines in multi-task manipulation w.r.t accuracy and robustness.
Related Works
Multi-task Learning in Manipulation
The core challenges for general-purpose robots lie in task generalization capabilities and multi-task execution in complex environments. In the past, multi-task learning [99, 100, 143, 101] has achieved remarkable progress. It primarily consists of two approaches: Modular solutions decompose skills into reusable basic units (e.g., perception, planning, and control modules) [95]. Though interpretable, they face the combinatorial explosion problem—as task complexity increases, the cost of designing interfaces and coordinating between modules grows exponentially, leading to system rigidity. For instance, ManipGen [98] consumed massive training resources to train over 1,000 task-specific experts. Conversely, end-to-end solutions employ a single dense model to directly map perception to action [38, 40, 41, 30, 42, 43, 97]. While simplifying the pipeline, significant differences in visual feature distributions and motion patterns across tasks (e.g., grasping fragile objects vs. tightening screws) cause the model to struggle with convergence during training due to parameter conflicts. The Mixture of Experts (MoE) [12] addresses the scaling bottleneck of large models through a sparse activation mechanism. SDP [14] integrated MoE into diffusion policies to address multi-task learning and task transfer. Differently, we propose a task-aware MoE (TaskMoE) to resolve feature conflicts between different robotic manipulation tasks, activating distinct task-specific experts for different tasks, thereby enhancing multi-task execution performance and generalization capabilities.
Vision-Language-Action Model
Recently, numerous VLA models [42, 38, 40, 143, 79] have been proposed. Unlike traditional approaches involving pose estimation [110, 112] and path planning [113, 114], this significantly streamlines the model pipeline. With the advancement of large language models (LLMs) and vision-language models (VLMs), these large models have been applied in robotics [39, 119, 73, 94, 74, 111], leveraging their powerful priors to provide world knowledge for robots, thereby enhancing model generalization and performance. However, most current VLA models rely on a single or fixed viewpoint, making it difficult for robots to achieve comprehensive perception of 3D space, which limits action prediction capabilities. To address this, we propose a perception scheme featuring dynamic multi-view rerendering to enhance the spatial perception and comprehension capabilities, consequently improving the robot’s execution performance in complex scenarios.
Figure 2: The overview of our Task-Aware View Planning (TAVP) framework. The input of this framework is four RGB-D images from fixed viewpoints. First, it converts them into point clouds and aggregates them into a global point cloud in the world coordinate system. Then, it diverges into two branches: One branch (orange) performs Coarse Grounding (a) to predict the approximate position of the end-effector. Subsequently, it moves the center of the global point cloud to this predicted position, performs scaling and cropping, retaining the important point cloud region. Another branch (green) receives the global point cloud, passes it through MVEP to predict the optimal camera parameters for the observation viewpoint. Then, using these parameters, it renders a 2D image from the point cloud processed by the red branch. This rendered image is fed into the Fine Grounding (b) to predict the final robot action, including the end-effector position, rotation, gripper status, and collision state.
Reinforcement Learning in Robotics
Reinforcement learning (RL) has achieved significant accomplishments in robotics [102]. However, implementing RL starting from scratch typically necessitates complex reward mechanism and training paradigm design [103, 108, 107, 105, 106]. Recently, several works [104, 88] have demonstrated promising results by using RL to fine-tune pre-trained models. To optimize our MVEPN (Multi-Viewpoint Exploration Policy Network), we designed pseudo-environment interactions and a view-aware reward function, enabling policy learning based on a pre-trained model using offline expert data. Furthermore, to adapt between the MVEPN and the action network, we introduce supervised post-training to coordinate the robot’s ability to “See and Act”, thereby enhancing its task execution performance.
Methodology
Overview
Our proposed Task-Aware View Planning (TAVP) framework aims to identify optimal viewpoints for accurate and robust robotic manipulation, guided by task-specific visual feature extraction. The full pipeline is illustrated in Fig. 2. TAVP takes as input a language instruction, the current visual observations from RGB-D cameras, and the current gripper state. To enable the model to explore the optimal observations from arbitrary viewpoints, we first reconstruct a 3D point cloud of the scene using the input RGB-D images. To narrow down the search space and enable task-aligned view selection, we leverage the coarse prediction stage from RVT-2 [143] to generate an area of interest.However, unlike RVT-2, which extracts visual features for all tasks using a shared Multi-View Transformer (MVT), we introduce a Task-Aware Mixture-of-Experts (TaskMoE) module before MVT to route instructions to specialized expert encoders. This design enables more precise and task-aligned visual feature extraction, which benefits both viewpoint selection and action prediction. Starting from the identified area of interest, we employ a Multi-Viewpoint Exploration Policy Network (MVEPN) to search for optimal camera poses that maximize the visibility of the target object and the end-effector. The selected viewpoints are then re-rendered into image observations and processed using another TaskMoE-based MVT before being passed to the action prediction model. For action prediction, we also upgrade the autoregressive action sequence model proposed in ARP [79] with our proposed TaskMoE to attain more task-specific feature extraction and action prediction.
In the following sections, we detail the design of: 1) the TaskMoE, 2) the MVEPN, 3) the training strategy.
TaskMoE
To address the inherent heterogeneity of complex manipulation tasks in multi-task learning, where different tasks often require substantially distinct visual representations and action policies, we introduce a task-aware Mixture-of-Experts module (TaskMoE), as illustrated in Fig. 3. Compared to prior MoE-based approaches [12, 14], our TaskMoE introduces two key innovations.
First, instead of relying solely on task identifiers for expert selection, we incorporate richer instruction- and scene-related cues to guide expert routing more effectively, which is crucial for accurate multi-task robotic manipulation. Specifically, as shown in Fig. 3, we design a cross-modality module that employs a cross-attention mechanism to model the interaction between instruction and visual information. The resulting context-aware features are then fused with the task identifier via a Feature-wise Linear Modulation (FiLM) layer [181], enabling more adaptive and task-sensitive expert selection.
Second, to improve the scalability and generalization of TaskMoE, we decouple the number of routing gates from the total number of tasks. Concretely, we allocate $N_{G}$ gates for all $N_{J}$ tasks, where $N_{G}<N_{J}$. This design not only accommodates task diversity but also promotes parameter sharing among tasks with similar visual or semantic characteristics. For example, as illustrated in Fig. 3, Task 1 and Task 2 (both involving opening drawer) are routed through the same gate but directed to different experts based on their specific operation requirements. In contrast, Task 3, which is semantically dissimilar, is routed through a different gate. This setup encourages the discovery of latent task clusters and provides the capacity to generalize to unseen tasks that share structural similarities with seen ones, thereby enhancing the transferability and robustness of TaskMoE. Notably, all tasks share a common pool of $N_{E}$ experts, and for each input, only the top-$k$ experts (based on gating scores) are activated to guide task-specific visual feature extraction.
Multi-Viewpoint Exploration Policy
Our proposed Multi-Viewpoint Exploration Policy (MVEP) aims to select $K$ viewpoints that maximally capture informative regions related to the manipulation target, thereby enhancing the accuracy of robotic action prediction. MVEPN takes as input the reconstructed point cloud $\mathcal{P}\in\mathbb{R}^{N\times 3}$ and its associated RGB features $\mathbf{F}_{\text{img}}\in\mathbb{R}^{N\times 3}$, which are concatenated to form the input representation:
$\mathbf{X}=\text{Concat}(\mathcal{P},\mathbf{F}_{\text{img}})\in\mathbb{R}^{N\times 6},$ (1)
where $N$ denotes the number of 3D points. The fused representation $\mathbf{X}$ is processed by a multi-layer perceptron (MLP), which predicts the parameters of $K$ camera poses.
In this work, we represent each camera pose using a look-at model [182], which decouples the camera position and orientation via spherical coordinates. Therefore, the explored views are parameterized as 5-dimensional vectors $\mathbf{p}^{i}=(\theta^{i},\phi^{i},r^{i},\theta_{\text{up}}^{i},\phi_{\text{up}}^{i})\in\mathbb{R}^{5}$, where $(\theta,\phi,r)$ specify the camera center in spherical coordinates relative to the origin, $(\theta_{\text{up}},\phi_{\text{up}})$ define the orientation of the up vector, and $i\in[0,K-1]$ means the $i_{th}$ view. So, all camera poses are $\mathbf{p}=\{\mathbf{p}^{i}|i=0,...,K-1\}$.
To facilitate gradient-based policy optimization using reinforcement learning, we model the camera pose parameters as samples from a Gaussian distribution. Instead of directly predicting the viewpoint $\mathbf{p}^{i}$, the network outputs the mean and log-standard deviation, enabling differentiable sampling via the reparameterization trick.
To facilitate gradient-based optimization of the viewpoint selection policy, we model each camera pose as a sample from a Gaussian distribution rather than predicting deterministic values. Specifically, for each of the $K$ viewpoints, our MVEPN outputs the mean and log-standard deviation of a diagonal Gaussian distribution over the 5-dimensional camera pose parameters. Formally, the output for the $i$-th viewpoint is denoted as $\mu^{i},\log\sigma^{i}]$, where:
$\mu^{i}$ $=[\mu_{\theta}^{i},\mu_{\phi}^{i},\mu_{r}^{i},\mu_{\theta_{\text{up}}}^{i},\mu_{\phi_{\text{up}}}^{i}],$ (2)
$\log\sigma^{i}$ $=[\log\sigma_{\theta}^{i},\log\sigma_{\phi}^{i},\log\sigma_{r}^{i},\log\sigma_{\theta_{\text{up}}}^{i},\log\sigma_{\phi_{\text{up}}}^{i}].$
Then the explored camera pose $\tilde{\mathbf{p}}^{i}=(\tilde{\theta}^{i},\tilde{\phi}^{i},\tilde{r}^{i},\tilde{\theta}_{\text{up}}^{i},\tilde{\phi}_{\text{up}}^{i})$ is sampled using the reparameterization trick:
$\tilde{\mathbf{p}}^{i}=\mu^{i}+\sigma^{i}\odot\epsilon^{i},\quad\epsilon^{i}\sim\mathcal{N}(\mathbf{0},\mathbf{I}),$ (3)
which enables end-to-end training via backpropagation.
To ensure the sampled poses remain valid in the spherical coordinate system, we constrain each component to lie within a normalized range using the sigmoid function:
$\tilde{\theta}^{i}$ $=\pi\cdot\sigma(\tilde{\theta}^{i}),\quad\tilde{\phi}^{i}=2\pi\cdot\sigma(\tilde{\phi}^{i}),$ (4)
$\tilde{r}^{i}$ $=r_{\min}+(r_{\max}-r_{\min})\cdot\sigma(\tilde{r}^{i}),$
$\tilde{\theta}_{\text{up}}^{i}$ $=\pi\cdot\sigma(\tilde{\theta}_{\text{up}}^{i}),\quad\tilde{\phi}_{\text{up}}^{i}=2\pi\cdot\sigma(\tilde{\phi}_{\text{up}}^{i}),$
where $\sigma(\cdot)$ denotes the sigmoid activation, and $r_{\min}$ and $r_{\max}$ define the allowable radial bounds.
Figure 3: Pipeline of the TaskMoE. Our proposed TaskMoE takes Task ID, Instruction, and Vision as inputs to guide expert selection for task-specific visual representation learning. To improve scalability and generalization, we design a compact gating mechanism with $N_{G}$ gates shared across $N_{J}$ tasks ($N_{G}<N_{J}$). This design allows tasks with similar action patterns (e.g., Task 1 and Task 2) to share the same gate, while assigning distinct gates to semantically diverse tasks (e.g., Task 3), thereby enabling effective feature specialization across a variety of manipulation tasks.
Training Strategy
The entire training process of our Task-Aware View Planning (TAVP) model consists of three stages:
Stage 1: In the first stage, we train a fixed-view variant of TAVP using three default rendering perspectives: front, left, and top. This stage follows a similar training pipeline to RVT-2 [143], where the overall loss comprises several components. Specifically, $\mathcal{L}_{hc}$ and $\mathcal{L}_{hf}$ denote the cross-entropy losses over heatmaps produced by the coarse and fine grounding modules, respectively. The ground-truth heatmaps are generated using a truncated Gaussian distribution centered at the 2D projection of the ground-truth 3D location, as described in RVT [155]. Additionally, the loss includes the end-effector rotation loss $\mathcal{L}_{rot}$, computed as the cross-entropy loss for each of the Euler angles, as well as the gripper state loss $\mathcal{L}_{gri}$ and collision indicator loss $\mathcal{L}_{col}$, both formulated as binary classification losses. The total objective for the first stage is given by:
$\mathcal{L}_{\text{s1}}=\mathcal{L}_{hc}+\mathcal{L}_{hf}+\mathcal{L}_{rot}+\mathcal{L}_{gri}+\mathcal{L}_{col}.$ (5)
Stage 2: In the second stage, we refine the Multi-View Exploration Policy (MVEP), using the fixed-view TAVP model trained in Stage 1 as a baseline. MVEP is optimized via the Proximal Policy Optimization (PPO) algorithm [93], which typically requires interaction with a physical environment to obtain reward feedback. To mitigate the high time cost of such interaction, we introduce a novel pseudo-environment interaction mechanism. The details of the mechanism are presented in the supplementary materials.
Specifically, we treat the fixed-view TAVP as a reference model to supervise the training of MVEP. Given the same observations and instructions, the reference model yields losses $\mathcal{L}_{\text{ref}}=[\mathcal{L}_{\text{hf}},\mathcal{L}_{\text{rot}},\mathcal{L}_{\text{gri}},\mathcal{L}_{\text{col}}]$. In contrast, MVEP generates dynamic viewpoints for action prediction, resulting in losses $\mathcal{L}_{\text{TAVP}}=[\mathcal{L}^{\prime}_{\text{hf}},\mathcal{L}^{\prime}_{\text{rot}},\mathcal{L}^{\prime}_{\text{gri}},\mathcal{L}^{\prime}_{\text{col}}]$. Taking the reference model as a performance lower bound, our objective is to make $\mathcal{L}_{\text{TAVP}}$ smaller than $\mathcal{L}_{\text{ref}}$. The first reward term is therefore defined as:
$r_{0}=\mathcal{L}_{\text{ref}}-\mathcal{L}_{\text{TAVP}}.$ (6)
In addition to the task loss reward, we incorporate a confidence-based reward derived from the fine grounding module. Specifically, we compute the negative average entropy across the grounding heatmaps:
$r_{1}=-\frac{1}{K}\sum_{i=1}^{K}\mathcal{H}(\text{softmax}(\mathbf{H}_{i})),$ (7)
where $\mathcal{H}$ denotes the entropy function and $\mathbf{H}_{i}$ is the heatmap for view $i$ from the fine grounding module.
To further encourage viewpoint diversity, we introduce an additional reward based on the average pairwise cosine distance between camera positions:
$r_{2}=\frac{1}{K(K-1)}\sum_{i\neq j}\left(1-\cos(\mathbf{p}_{i},\mathbf{p}_{j})\right),$ (8)
where $\mathbf{p}_{i}$ represents the camera position for view $i$.
The overall reward is computed by adaptively normalizing and aggregating all components:
$r=\sum_{i=0}^{2}w_{i}\cdot\mathcal{N}(r_{i}),$ (9)
where $w_{i}$ are learnable component weights and $\mathcal{N}$ is an online normalization operator. The adaptive normalization maintains the mean and variance of each component using the Welford algorithm [121] for normalization, and clips the final reward to the range [10,10][-10,10] to ensure training stability.
In this stage, only MVEP is trainable, and all the other components of TAVP are frozen.
Stage 3: In the third stage, to enable MVEP to better adapt to action generation, we further fine-tune the entire TAVP model, excluding MVEP, using the same loss functions as in Stage 1.
Method Avg. Success Close Jar Drag Stick Insert Peg Meat off Grill Open Drawer Place Cups Place Wine Push Buttons
RVT2 81.4 100.0 99.0 40.0 99.0 74.0 38.0 95.0 100.0
ARP 81.6 97.6 88.0 53.2 96.0 90.4 48 92.0 100.0
ARP+ 84.9 95.2 99.2 78.4 97.6 92.8 48.8 96.0 100.0
TAVP(Ours) 86.7 100.0 100.0 96.0 100.0 88.0 44.0 96.0 100.0
Method Put in Cupboard Put in Drawer Put in Safe Screw Bulb Slide Block Sort Shape Stack Blocks Stack Cups Sweep to Dustpan Turn Tap
RVT2 66.0 96.0 96.0 88.0 92.0 35.0 80.0 69.0 100.0 99.0
ARP 68.0 99.2 94.4 85.6 98.4 35.2 55.2 76.8 90.4 100.0
ARP+ 69.6 98.4 86.4 89.6 92.8 46.4 63.2 80.0 97.6 96.0
TAVP(Ours) 76.0 100.0 92.0 88.0 100.0 64.0 68.0 60.0 92.0 96.0
Table 1: Performance comparison across 18 tasks in RLBench. We compare the success rate of each individual task and the average success rate across all tasks between dense fixed-view baselines such as RVT2, ARP, and ARP+ and our TAVP.
Experiments
To fully evaluate our Task-aware View Planning (TAVP) model, we conducted comprehensive experiments in both simulated and real-world environments. Detailed experimental settings for the real-world and the corresponding results are presented in the supplementary materials.
Experimental Setup
Simulation Setup.
Our simulation experiments are conducted on RLBench dataset [19], a multi-task learning benchmark implemented in the CoppeliaSim simulator featuring a 7-degree-of-freedom (7-DoF) Franka Emika Panda robot in a tabletop environment. We employ a curated subset comprising 18 distinct manipulation tasks, each containing multiple variations (tasks details can be found in the supplementary materials). Observations consist of RGB-D images ($128\times 128$) captured from four fixed views: frontal, left shoulder, right shoulder, and wrist-mounted. The action space represents the end-effector pose as $\mathbf{a}=(\mathbf{p}_{\text{xyz}},\mathbf{q}_{\text{rot}},g_{\text{gripper}},c_{\text{collision}})\in\mathbb{R}^{7}\times\{0,1\}^{2}$, where $\mathbf{p}_{\text{xyz}}\in\mathbb{R}^{3}$ denotes Cartesian coordinates, $\mathbf{q}_{\text{rot}}\in\mathbb{R}^{4}$ is the quaternion rotation, $g_{\text{gripper}}\in\{0,1\}$ controls gripper state, and $c_{\text{collision}}\in\{0,1\}$ indicates collision.
Implementation Settings.
In our experiments, the number of camera viewpoints $K=3$, the number of gates in TaskMoE $N_{G}=8$, and the number of task experts $N_{E}=16$. For each task, the Top2Top-2 task experts are selected. The default radial constraint for the camera is $r_{\text{min}}=0.75$ m and $r_{\text{max}}=1.3$ m. Our TAVP model was trained using $4\times$ NVIDIA RTX A800 GPUs with 80GB memory. During testing, we sequentially validated each task using $1\times$ NVIDIA RTX A800 GPU and computed the average success rate.
Main Results
Results Across 18 RLBench Tasks. Compared to dense models that employ fixed camera viewpoints for re-rendering, our sparse Task-Aware View Planning (TAVP) model demonstrates significant performance gains in multi-task scenarios. We conduct benchmark evaluations on 18 distinct RLBench tasks, encompassing a diverse range of manipulation challenges. As shown in Table  1, RVT2, ARP, and ARP+ (where ARP and ARP+ differ only in model capacity) are all dense models utilizing fixed viewpoints for re-rendering. In contrast, our model employs the sparse TAVP approach. The results indicate that our model achieves substantially higher success rates across all tasks compared to the strongest dense baseline, ARP+.
Moreover, the per-task analysis reveals that TAVP consistently outperforms dense baselines in the tasks with the susceptibility to occlusion or fine-grained perceptual demands. In these tasks (e.g., put in cupboard and insert peg), TAVP achieves up to a +56% improvement by repositioning cameras to enhance the joint visibility of the target and end-effector. It indicates that the robot’s gripper or neighbouring objects routinely block the main line of sight, and dynamic views are essential for success. Even on seemingly less challenging operations (close jar, meat off grill, and slide block), TAVP registers increments of 4%, indicating that its benefits are not confined to a handful of corner cases. Averaged over all tasks, the success rate rises from 84.9% (ARP+) to 86.7%, confirming that the combination of active view planning and task-aware feature routing delivers domain-wide gains rather than isolated spikes in performance.
Configuration Average Success Rate (%)
TAVP 86.67
w/o TaskMoE 85.56
w/o Active Exploration Fine-tuning
(Random Viewpoints)
8.89
Table 2: Ablation study results in RLBench tasks.
$K$ $r$ PID $T$ PGC PMS CJ SR
[…]
Table 3: Performance under different view numbers $K$ and camera radial constraints $r$, which $\Delta(r_{min},r_{max})=\pm 20\%$. Additionally, task PID $\to$ Put item in drawer, $T$ $\to$ Turn tap, PGC $\to$ Put groceries in cupboard, PMS $\to$ Put money in safe, CJ $\to$ Close jar.
Ablation Studies.
To verify the effectiveness of each module in our TAVP, we conducted comprehensive ablation studies on 18 RLBench tasks, as shown in Table 2. The full TAVP configuration achieves an 86.44% average success rate across all tasks. Removing the TaskMoE architecture from both feature extraction and action generation components (w/o TaskMoE) causes a moderate performance degradation to 85.56%, indicating its importance in multi-task representation learning. Crucially, replacing our actively explored viewpoints with random viewpoints (w/o Active Exploration) yields only 8.89% success, demonstrating that task-aware view planning contributes dominantly to the overall performance gains.
Method Average Success Rate (%) Average Inference Time(s)
ARP+ 84.90 0.394
TAVP 86.67 0.436
Table 4: Comparison of the average success rate and inference time between ARP+ and TAVP across 18 tasks.
TaskMoE PID $T$ PGC PMS CJ SR
Open drawer
(New task)
32.0 92.0 16.0 32.0 76.0 49.6 12.0
$\times$ 0.0 84.0 0.0 12.0 24.0 24.0 0.0
Table 5: Under the settings of $K=3$, $r_{\text{min}}=0.75$, and $r_{\text{max}}=1.3$, a comparison of performance on in-domain tasks (ID) and out-of-domain unseen tasks (OOD) with and without TaskMoE. Additionally, task PID $\to$ Put item in drawer, $T$ $\to$ Turn tap, PGC $\to$ Put groceries in cupboard, PMS $\to$ Put money in safe, CJ $\to$ Close jar.
Quantitative Analysis
To comprehensively evaluate our TAVP model, we conduct extensive experiments on five representative robotic manipulation tasks, as shown in Table 3. Our analysis encompasses three critical dimensions: 1) sensitivity analysis examining key model parameters including view quantity and camera radial constraints ($r_{min}$$r_{max}$), 2) performance-efficiency trade-off analysis quantifying inference latency against task success metrics, and 3) zero-shot generalization analysis of the TaskMoE component on unseen tasks.
Validation Analysis Experimental Setup. The validation analysis experiments were conducted on the RLBench dataset. We selected five manipulation tasks: put item in drawer, turn tap, put groceries in cupboard, put money in safe, and close jar. During training, each task utilized 20 demonstration episodes. For testing, each task was evaluated over 25 episodes, with the success rate recorded for each individual task and the average success rate calculated across all tasks. In the zero-shot generalization analysis experiment, the unseen task “Open drawer” was selected for evaluation. Similarly, this task was tested over 25 episodes. For the performance-efficiency trade-off analysis experiment, we maintained the setup described in Simulation Setup Simulation Setup..
Figure 4: Visualization of our TAVP and baseline model in simulation RLBench Environment and Real-world Environment.
Sensitivity Analysis.
We conducted ablation studies to analyze the key parameters of our TAVP model, including the camera radial constraint and the number of views. The radial constraint was evaluated in three cases based on a baseline range of ($0.75\sim 1.3$)m, which was varied by $\pm 20\%$. The experimental results, as shown in Table 3, indicate that compared to the baseline range, a smaller and tighter constraint ($0.90\sim 1.04$)m improved the model’s performance across all tasks, with the average success rate increasing significantly from $49.6\%$ to $56.0\%$. Conversely, a larger and looser constraint ($0.60\sim 1.56$)m led to a slight decrease in performance, with the average success rate dropping to $48.8\%$. This suggests that providing the model with a more precise and focused range of advantageous viewing distances can effectively enhance its final task performance.
In addition to radial constraints, we investigated the impact of varying the number of camera views on task performance. As shown in Table 3, under the same radial constraint range of $(0.75\sim 1.3)\text{m}$, increasing the number of views from 2 to 4 consistently improved performance across all five tasks. Specifically, the average success rate rose from $27.2\%$ (2 views) to $49.6\%$ (3 views), and further to $55.2\%$ with 4 views. Notably, performance gains were especially prominent in occlusion-prone tasks such as ”Put money in safe”, indicating that multi-view observations significantly enhance spatial perception and reduce failure caused by partial observability, enabling the TAVP to generalize better and make more informed decisions during manipulation tasks.
Performance-Efficiency Trade-off Analysis.
During inference, TAVP must predict dynamic multi-view camera poses and continuously adjust the rendering camera pose, resulting in a non-negligible computational overhead. To assess the practicality of the proposed TAVP model, we compare its average task success rate and inference latency with those of the baseline model, ARP+. The results, summarized in Table 4, show that TAVP achieves a higher average success rate than ARP+, while the increase in inference latency remains relatively modest (approximately 10.7%). These findings suggest that TAVP effectively balances task performance and operational efficiency, with design elements such as sampling acceleration and camera caching playing a key role in enhancing model efficiency.
Zero-shot Generalization Analysis.
We experimentally validated the generality and generalization capability of the TaskMoE module in the TAVP model, with the results presented in the Table 5. The experimental results indicate that, while keeping other parameters identical and at their default values, the inclusion of the TaskMoE module significantly enhances the model’s overall performance and generalization capability. On tasks seen during training, the model equipped with the TaskMoE module (✓) achieved an average success rate of 49.6%, more than double the 24.0% rate of the model without it ($\times$). This shows that the TaskMoE module enables more effective learning and execution of diverse, known tasks. For new, unseen tasks (OOD, Out-of-Distribution), the model without the TaskMoE module failed completely, with a success rate of 0. In contrast, the model with the TaskMoE module achieved a 12.0% success rate on the new task “Open drawer”, demonstrating a performance breakthrough. This proves that the TaskMoE endows the model with the ability to generalize learned knowledge to new tasks. In conclusion, the TaskMoE module not only enhances the model’s performance on known tasks but, more crucially, improves its generalization capability for new tasks, thereby validating the effectiveness of its design.
Visualization Results.
To validate the effectiveness of our method, we demonstrate the performance of TAVP and ARP+ in both the simulation environment RLBench and the real world. Figure 4 highlights TAVP’s advantage over the fixed-view ARP+ across simulation and reality. In RLBench, ARP+ loses sight of the mug once it is lifted, cannot perceive the cupboard interior, and drops the coffee mid-air; TAVP re-renders three task-guided views that simultaneously capture the mug and cupboard, yielding a correct 6-DoF placement. On the real robot, ARP+ misjudges the plate’s height, opening the gripper too early so grapes roll off target, whereas TAVP selects a clarifying side view and a persistent wrist view, preserving depth cues and completing the drop cleanly. By eliminating occlusion and recovering spatial context, TAVP turns visual completeness directly into manipulation success, validating that dynamic “seeing” underpins robust “acting.”
Conclusion
Our TAVP framework establishes that dynamic view planning and task-aware representation learning significantly advance robotic manipulation capabilities. The MVEP module’s viewpoint optimization effectively overcomes occlusion limitations in fixed-viewpoint systems, while TaskMoE’s specialized feature extraction mitigates multi-task interference. These innovations collectively enhance performance across diverse manipulation challenges and enable meaningful generalization to novel tasks.
Appendix
In this supplementary material, we provide additional details and results of our proposed TAVP to complement the main paper. The content is organized as follows:
  • Appendix A: Real-world experimental details and results. Corresponding demo videos of the results are also included in the supplementary materials for a more intuitive understanding.
  • Appendix B: Additional design details of TAVP, including point cloud aggregation, camera pose parameterization, and the proposed pseudo-environment interaction mechanism.
  • Appendix C: Implementation details and information about the RLBench benchmark. Besides, we also provide more visualized results of RLBench [19], whose corresponding demo videos also included in the supplementary materials.
  • Appendix D: Discuss and analyze the limitations of our TAVP model.
Appendix A. Real-World Experimental Details and Results
Real-robot Setup.
We employ the Dobot Xtrainer platform featuring two 6-degree-of-freedom (6-DoF) Nova 2 collaborative manipulators, equipped with three Intel RealSense D405 depth cameras mounted at fixed positions: one at an overhead lateral position and two wrist-mounted cameras on both manipulators, as shown in Fig 5. Five distinct manipulation tasks were designed for evaluation the effectiveness of our proposed TAVP, including arrange_tableware, boil_water, cover_lid, pick_grape, and press_button. Expert demonstrations were collected, yielding 100 demonstration episodes for every task for training. Some samples are provided in Fig. 6. To accelerate the efficiency of training, we extract key frames from the collected demos and only retain important key frame data.
Real-world Experimental Results.
Fig. 7 shows the real-world experimental results. In five real-world tasks, our TAVP performs well in tasks such as boil_water, cover_lid, pick_grape, and press_button. Especially in the pick_grape task, it successfully executed despite occlusion of the manipulation target in one viewpoint, demonstrating the advantage of our model in global information perception. Corresponding demo videos are also provided for a more intuitive understanding of the real-world results.
Figure 5: Real-World Environment Setup.
Figure 6: Demonstrations of 5 real-world tasks and their corresponding language instructions.
Figure 7: Visualization of Robots’ Execution Processes in 5 Real-World Environments. In the figure, the T-axis represents the time axis, indicating the duration from the start to the end of the task. Each demonstration showcases views from two camera perspectives—the side camera and the wrist camera—along with the language instruction.
Appendix A Appendix B. More design details of TAVP
Point Cloud Aggregation.
In the RLBench dataset experiments, we select RGB-D images from four fixed viewpoints: front, right_shoulder, left_shoulder, and wrist as inputs. Following the approach of RVT-2 [143], we convert the depth maps from all viewpoints into point clouds in a global coordinate system using the intrinsic and extrinsic camera parameters. Subsequently, we perform multi-view point cloud aggregation and remove redundant points outside the workspace to obtain the final scene point cloud, as illustrated in Fig. 8.
Figure 8: The aggregated full-scene point cloud of the put_groceries_in_cupboard task from multiple viewpoints
Pseudo-environment Interaction Mechanism.
To enable efficient policy optimization without physical environment interaction, we design a novel pseudo-environment interaction mechanism. As shown in Algorithm 1. The shadow network $\pi_{\text{shadow}}$ (a frozen copy of the policy before training) processes observation $\mathbf{o}$ to compute reference losses $\mathcal{L}_{\text{ref}}$. MVEP $\pi_{\theta}$ processes $\mathbf{o}$ to generate camera poses $\mathbf{p}$ ($\mathbf{p}$ as policy action), the rendered observation images from multi-view camera poses are fed into TaskMoE and TaskMoE-ARP (Autoregressive Action Policy) to compute the current policy losses $\mathcal{L}_{\text{TAVP}}$. The reward $r$ is then calculated based on $\mathcal{L}_{\text{ref}}$, $\mathcal{L}_{\text{TAVP}}$, and $\mathbf{p}$. Finally, the experience tuple $(\mathbf{o},r,\mathbf{p},\log\pi_{\theta_{\text{old}}}(\mathbf{p}|\mathbf{o}),V_{\theta_{\text{old}}}(\mathbf{o}))$ is stored in the replay buffer $\mathcal{B}$ for policy updates. This approach enables batch environment interaction, leveraging existing demonstration data to improve policies and enhance data utilization efficiency.
Algorithm 1 Pseudo-Environment Interaction
1:Input: Current observation $\mathbf{o}$, shadow network $\pi_{\text{shadow}}$
2:Output: Transition tuple $\tau$
3:# Compute reference loss
[…]
5:# Generate camera poses & compute current model loss using TAVP
[…]
7:# Compute reward
[…]
9:# Construct transition tuple
10:return $(\mathbf{o},\mathbf{p},r,\log\pi_{\theta_{\text{old}}}(\mathbf{p}|\mathbf{o}),V_{\theta_{\text{old}}}(\mathbf{o}))$
Camera Pose Parameterization
The camera pose is mathematically represented using a look-at model defined by a 5-dimensional parameter vector $\mathbf{p}^{i}=(\theta^{i},\phi^{i},r^{i},\theta_{\text{up}}^{i},\phi_{\text{up}}^{i})\in\mathbb{R}^{5}$($i\in[0,K-1]$, $K$ is number of camera viewpoint). This representation decouples camera position from orientation through spherical coordinates:
$\mathbf{t^{i}}$ $=\begin{bmatrix}x^{i}\\ y^{i}\\ z^{i}\end{bmatrix}=\begin{bmatrix}r^{i}\sin\theta^{i}\cos\phi^{i}\\ r^{i}\sin\theta^{i}\sin\phi^{i}\\ r^{i}\cos\theta^{i}\end{bmatrix}$ (10)
where $\mathbf{t}^{i}$ denotes the camera center position. The viewing direction is intrinsically defined as $\mathbf{v}_{\text{look}}^{i}=-\mathbf{t}^{i}/\|\mathbf{t}^{i}\|_{2}$ toward the origin. The up-vector orientation is parameterized as:
$\mathbf{v}_{\text{up}}^{i}=\begin{bmatrix}\sin\theta_{\text{up}}^{i}\cos\phi_{\text{up}}^{i}\\ \sin\theta_{\text{up}}^{i}\sin\phi_{\text{up}}^{i}\\ \cos\theta_{\text{up}}^{i}\end{bmatrix}$ (11)
The final camera matrix is constructed via Gram-Schmidt orthogonalization between $\mathbf{v}_{\text{look}}^{i}$ and $\mathbf{v}_{\text{up}}^{i}$.
Figure 9: Demonstrations of 18 RLBench tasks and their corresponding language instructions.
Appendix B Appendix C. More Experimental Details and Resutls
Implementation Details.
In the RLBench experiments, during the supervised pre-training phase, our experimental hyperparameter settings comply with APR[79]. In the offline reinforcement learning phase for MVEP, our hyperparameter configurations are detailed in Table 6.
Hyperparameter Value
MVEP
$K$ number of the camera viewpoints 3
$N_{MVEP}$ The number of input point clouds for MVEP 2048
embedding size 512
sensitive parameters
$\theta$ $0\sim\pi$
$\phi$ $0\sim 2\pi$
$r_{min}$ 0.75
$r_{max}$ 1.3
Train & Eval
observation RGBD $4\times 128\times 128\times 4$
re-render image resolution $224\times 224$
maximum evaluation steps 25
train epochs 20
eval frequency 100
batch size 96
learning rate 2.0e-6
learning rate scheduler cosine
optimizer LAMB
Table 6: Hyperparameters for offline reinforcement learning training in RLBench experiments.
More Details of RLBench.
To evaluate the effectiveness and generalization capability of the proposed TAVP for multi-task robotic manipulation, we conduct experiments on 18 diverse tasks from the RLBench benchmark, including close jar, stack blocks, and so on. TAVP is trained with 100 demonstrations per task. Example demonstration samples for each task are shown in Fig. 9. During evaluation, the model is tested with 25 demonstrations using the same settings. For a more intuitive understanding, we also include corresponding video results.
Appendix C Appendix D. Discussions and Limitations.
Our TAVP framework establishes that dynamic view planning and task-aware representation learning significantly advance robotic manipulation capabilities. The MVEP module’s viewpoint optimization effectively overcomes occlusion limitations in fixed-viewpoint systems, while TaskMoE’s specialized feature extraction mitigates multi-task interference. These innovations collectively enhance performance across diverse manipulation challenges and enable meaningful generalization to novel tasks.
However, the framework has certain limitations: there is a slight increase in inference latency due to active view planning; it relies on accurate global point clouds, making it difficult to handle reflective/transparent objects in real-world environments. Future work could explore multi-sensor integration and domain adaptation techniques to enhance real-world robustness. The modular design of TAVP components offers promising pathways for extending these capabilities to broader manipulation scenarios.